NEAR SYMMETRY IN OUR UNIVERSE: THE ORIGIN OF LIFE AND THE PREPONDERANCE OF MAT-TER OVER ANTIMATTER

\underline{P. R. BUNKER}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The repercussions of the slight breakdown of some of the symmetries in nature will be discussed with regard to their affect on the origin of life and on the preponderance of matter over antimatter in our Universe. It will be explained how ultra high resolution molecular spectroscopy experiments coupled with ultra high level ab initio quantum chemical calculations may play a crucial role in furthering our understanding of these matters.